OpenStudy (anonymous):

A linear function has a slope of -7/9 and a y-intercept of 3. How does this function compare to the linear function that is represented by the equation y+11=-7/9(x-18)? A. It has the same slope and the same y-intercept. B. It has the same slope and a different y-intercept. C. It has the same y-intercept and a different slope. D. It has a different slope and a different y-intercept.

OpenStudy (anonymous):

It is A cause y+11=-7/9(x-18) you do distributive property so -7/9x+14 and you subtract 11 on both side to get y itself so y=-7/9+14 -11 y=-7/9x+3 so it stay the same

OpenStudy (anonymous):

one more?

OpenStudy (anonymous):

what is the y-intercept of y=2x-2